what is the formula computing the interest rate on an investment?
Answer #1

Assume an investment of I at the beginning of the period.

Let r be the rate of interest per compounding period and N be the number of years

Then, the amount A at the end of the compounding period after N years

A = I *(1+r)^N

Where Compond Interest , CI = A - I

Interest Rate on the investment , r = (A / I)^1/N -1